OpenStudy (anonymous):

Solve the equation for Y: 4x-10y=12

OpenStudy (anonymous):

Add '10y' to each side of the equation 4x + -10y + 10y = 12 + 10y Combine like terms: -10y + 10y = 0 4x + 0 = 12 + 10y 4x = 12 + 10y Divide each side by '4'. x = 3 + 2.5y Simplifying x = 3 + 2.5y